OpenStudy (anonymous):

Could someone help with this? I have no idea how to do it. You want to estimate the difference in grade point averages between two groups of university students to be accurate within 0.2 grade point, with probability approximately equal to 0.95. If thestandard deviation of the grade point measurements is approximately equal to 0.6, how many studentsmust be included in each group? (Assume that the groups will be of equal size)

OpenStudy (anonymous):

i tried this question but i get 0.029 as my answer

OpenStudy (kropot72):

If the difference is to be within 0.2 grade point, then each grade point average must be accurate to within 0.1 grade point, with the required approximate probability. The margin of error must be no more than 0.1 grade point for each grade point average. The margin of error is defined as: \[\Large z.\sigma _{\bar {x}}=z\frac{\sigma}{\sqrt{n}}\] where z is the critical value of z, which is z = 1.96 in this case, and n is the sample size. Plugging in values we get: \[\large 0.1=1.96\times\frac{0.6}{\sqrt{n}}\ ..........(1)\] Now all you need to do is to solve equation (1) to find the value of n, the sample size.

OpenStudy (anonymous):

nvm i got it. Sorry i didn't tell you. I got 35 as my answer. Could you help with another problem ?

OpenStudy (kropot72):

How did you solve it? My calculation gives a very different answer from yours. Please post the other problem as a new question.